40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 94     Entry time: Sun Aug 21 08:51:22 2016
Author: Gabriele 
Type: General 
Category: Daily Progress 
Subject: New disk installed 

Vented the chamber. Installed a new disk (MO 03). The one I measured yesterday is now named MO 02 and it is the one with the "burnt mark" from the previous experiment (due to the electrostatic drive).

Startep roughing pump at 8:50am. Started turbo pump at 9.00am.

Excited the disk at 9:50:30am with white noise, amplitude 10 V. Pumps are still running, pressure is about 2e-6 Torr

At 11:10am I stopped the roughing pump, pressure is 1.4e-6 Torr. Exciting again the disk at 11:10:45am. At 12:25am I checked again the situation, since both measurements look quite weird, especially for the first couple of modes.

I think the reflection from the balck glass is interfering with the reflection from the disk. Probably I wasn't careful enough when I aligned the disk. At about 12:35 I stopped the turbo pump. I'm going to open the chamber and realign everything again.

ELOG V3.1.3-